OpenStudy (macaylam):

Please help! (medal + fan) Choose the expressions that are sums or differences of two cubes. 64 + a^12b^51 –t^6 + u^3v^21 8h^45 – k^15 75 – n^3p^6 –27 – xz^9

OpenStudy (sunnnystrong):

Recall that: Sum of cubes: \[a^3 + b^3 = (a + b)(a^2 – ab + b^2) \] Difference of cubes: \[a^3 – b^3 = (a – b)(a^2 + ab + b^2)\]